The spacecraft continues to be in excellent health,
and is now about 16 million km from Earth. Key activities completed this
week include successful completion of the K=15, R=1/6 convolutional code
tests and resolution of the attitude control software glitch detected last
week. Attitude control software has been re-enabled and is currently operating
nominally. In addition, we verified that the noise seen during ASI/MET health checks
is due to the Propulsion Drive Electronics. This noise appears to be radiative
in nature, and will not be an issue for surface operations because the PDE
is located on the cruise stage.
The Uplink Problem Tiger team has developed a plausible explanation for the majority of the command rejections and the CDU In Lock conditions. It involves harmonics from the uplink sweep locking up the CDU and pulling it away from the nominal command frequency. The team is developing a test plan to confirm this hypothesis and is also gathering information about the incidents where the CDU went into lock while we were not tracking.
An Operational Readiness Test (ORT) of the Sol 1-2 sequences was run on Jan 27 and 28. The sequences used were identical to the last pre-launch surface ORT. The ORT was successful in that all of the sequences were executed properly by the simulated lander and rover. However, a number of relatively minor problems were logged during the test. These problems were reviewed and action has been assigned in all cases for problem resolution.
Nineteen investigators have been selected by NASA Headquarters in response to the Announcement of Opportunity for selection of Mars Pathfinder Participating Scientists and a Facility Instrument Science Team for the Atmospheric Structure Instrument/Meteorology Package. An "All Hands" Pathfinder Science Team meeting has been set up for Feb. 5-7, 1997 at JPL to begin integrating the new investigators into the Experiment Operations Team.
Back to the Mars Pathfinder Home Page